I have no clue where to begin with this problem:
An investment firm is considering two alternative investments, A and B, under two possible future sets of economic conditions good and poor. There is a .60 probability of good economic conditions occurring and a .40 probability of poor economic conditions occurring. The expected gains and
losses under each economic type of conditions are shown in the following table:
Investment
Economic Conditions
Good Poor
A $380,000 -$100,000
B $130,000 $85,000
Using the expected value of each investment alternative, determine which should be selected.

Answer by Boreal(15235)

Expected values= Sum of X*p(x)
A: The plus is 380000*.6-100000(.4)=228000-40000=$188,000
B. 130000*0.6+85000(0.4)=78000+34000=$112,000
A has a better expected value so mathematically it should be selected.